During laparoscopic surgery, carbon dioxide gas is used to expand the abdomen to help create a larger working space.

If 5.00 L of CO2 gas at 19 ∘C at 785 mmHg is used, what is the final volume, in liters, of the gas at 31 ∘C and a pressure of 753 mmHg , if the amount of CO2 remains the same?
Answer #1

Ideal gas, pr=nRT p. Presur x-volume n=male of gas At constant n, (amount s ame) T = temp (K) 3) c = 273+ 31 K= 304 k Igc = 2